1. A local Little League plans to invest $10,000 to host a tournament. They expect to sell $15,000 worth of tickets and concessions that weekend, but if it rains, they won’t sell any tickets and will lose all of their money. The weather forecast for the weekend has a 20% chance of rain. What is the expected net value of the tournament?

$2,000

$3,000

$7,000

$12,000

2. Maurice Allais noticed that people’s tolerance for risk changes according to the situation. Given the set of choices below, answer the question that follows.

Choose gamble A or B
Gamble A Gamble B
A lottery ticket that pays $500,000 1% of the time, $50,000 95% of the time, and $0 4% of the time No gamble—receive $50,000 in cash 100% of the time
Choose gamble C or D
Gamble C Gamble D
A lottery ticket that pays $500,000 5% of the time A lottery ticket that pays $250,000 6% of the time



Which of the following combinations should a risk-neutral person select?

A and C

A and D

B and C

B and D

3. One game gives you a choice of receiving a guaranteed $900 or a 90% chance of winning $1,000. Another game has offered you $1,000 if you’re willing to take a 10% chance of losing it all. The utility associated between the two gambles may be different. This can be explained by which concept?

framing

prospect theory

preference reversal

status quo bias
Answer #1

1) Expected sales of the tickets and concessions = 0.2* 0 + 0.8*15000 = $ 12000

So the expected net value of the tournament = Expected sales - investment = 12000 - 10000 =$2000